30. Solve: r-6 X-4 + 16 = 0. Show work. r? -16

PLEASE WRITE THE ANSWER LEGIBLE!
Answer #1

30 + n-6 Sol 16 2²-16 2-4 (2-6)(x+4) +16 x² - 16 22 -6+40-24 + 16 = 0 ? - 2X-8 =0 tox? -4%+201-8=0 x (x-4)+2(x-4) 20 (x + 2)